# {refsection}

### v2

@@ C.   Fragmentation and aggregation experiments
The fragmentation and aggregation experiments involved subjecting a suspension of Microcystis colonies to the cone-and-plate shear and using brightfield microscopy to image the colonies under shear. The lower surface was positioned on top of an inverted microscope Nikon TS100 with a 4x objective lens. The center of the field of view was located $15~{}mm$ from the rotation axis and with the focal plane set $0.5~{}mm$ above the lower surface. Illumination was provided by a Schott KL 1500 HAL light source equipped with a cyan filter (low pass with a cut-off wavelength of 555 $nm$) to minimize the photosynthetic activity of the sample. Images were captured with a Prime BSI Express sCMOS camera with a resolution of $1.63~{}\mu m$ per pixel. Prior to each experiment, a reference background image was created using distilled water, which was then discarded. A volume of $3.5~{}ml$ of the suspension of Microcystis colonies was pipetted in between the shear surfaces. The shear protocol was as follows: pre-shear at $\dot{\varepsilon}=1.1$ ${m^{2}/s^{3}}$ for 1 minute to mix the suspension; 12 shear cycles of 5 minutes each followed by 12 cycles of 20 minutes each at the target dissipation rate (imaging cycles, at moderate dissipation rate of $0.019$ ${m^{2}/s^{3}}$ for 1 min, were intercalated with each shear cycle). During the imaging cycles, 200 frames were captured at a rate of 10 $fps$, ensuring that each frame was sampled over a different set of colonies. Examples of the captured frames are shown in Figure 1.
@@ Acknowledgments
  • Small colonies ($l<5$): The frame is cropped into a smaller section (400x400 pixels), to speed the processing. The background is subtracted using the reference image. The image intensity is renormalized and then thresholded. Small features ($50.5<l<$ 5) are segmented, measured, and counted.
@@
The turbulent shear was generated by a cone-and-plate shear, with the dimensions depicted in Fig. S7A. The flow regime can be identified from the average energy dissipation rate $\dot{\varepsilon}$ curve against the angular velocity $\omega$, with a $\sim\omega^{2}$ dependence in the laminar regime and a $\sim\omega^{5/2}$ dependence in the inertial region (Fig. S7B) [sdougos1984secondary, grad2000spectral]. The Kolmogorov scale $\eta$, which is an estimate of the size of the smallest eddies in a turbulent flow, can be calculated from the relation $\eta=\nu^{3/4}\dot{\varepsilon}^{-1/4}$. The Kolmogorov scale in our results range from $\eta\sim 85~{}\mu$m for the lowest dissipation rate ($\dot{\varepsilon}=0.019~{}{m^{2}/s^{3}}$) to $\eta\sim 18~{}\mu$m for the highest dissipation rate ($\dot{\varepsilon}=9.6~{}{m^{2}/s^{3}}$).
@@
[…]
@@
Fig. S8: Calibration steps for computing the correction function for the biovolume distribution. (A-B) Uncorrected concentration distribution of colonies measured in the cone-and-plate shear setup. Total number of counted colonies, $N=10776$. (C-D) Reference distribution measured in the inverted microscope after sedimentation of the colonies. Total number of counted colonies in panels C and D: $N=3066$ and $1455$, respectively. (E-F) Ratio between the reference and uncorrected concentration distributions Dashed lines indicate the fitted calibration parameters. (G-H) Normalized biovolume distribution measured in the cone-and-plate shear setup after the correction function. Error bars indicate the root mean squared of counting uncertainty and the size measurement uncertainty. Panels A,C,E,G display the distributions for small colonies ($l<5$), while panels B,D,F,H display the distributions for large colonies ($l>5$).

### v3

@@ C.  Fragmentation and aggregation experiments
The fragmentation and aggregation experiments involved subjecting a suspension of Microcystis colonies to the cone-and-plate shear and using brightfield microscopy to image the colonies under shear. The lower surface was positioned on top of an inverted microscope Nikon TS100 with a 4x objective lens. The center of the field of view was located 15mm15~$m$ from the rotation axis and with the focal plane set 0.5mm0.5~$m$ above the lower surface. Illumination was provided by a Schott KL 1500 HAL light source equipped with a cyan filter (low pass with a cut-off wavelength of 555 $nm$) to minimize the photosynthetic activity of the sample. Images were captured with a Prime BSI Express sCMOS camera with a resolution of $1.63~\mu m$ per pixel. Prior to each experiment, a reference background image was created using distilled water, which was then discarded. A volume of 3.5ml3.5~ml of the suspension of Microcystis colonies was pipetted in between the shear surfaces. The shear protocol was as follows: pre-shear at $\dot{\varepsilon}=1.1$ $m^{2}/s^{3}$} for 1 minute to mix the suspension; 12 shear cycles of 5 minutes each followed by 12 cycles of 20 minutes each at the target dissipation rate (imaging cycles, at moderate dissipation rate of $0.019$ $m^{2}/s^{3}$} for 1 min, were intercalated with each shear cycle). During the imaging cycles, 200 frames were captured at a rate of 10 fpsfps, ensuring that each frame was sampled over a different set of colonies. Examples of the captured frames are shown in Figure 1.
@@ Acknowledgments
  • Small colonies ($l<5$): The frame is cropped into a smaller section (400x400 pixels), to speed the processing. The background is subtracted using the reference image. The image intensity is renormalized and then thresholded. Small features ($0.5<l<5$) are segmented, measured, and counted.
@@
Fig. S5: Rheology of concentrated colonies of Microcystis strain V163. (A) Shear stress as a function of the shear rate obtained for a steady shear test. Solid line indicates a Herschel–Bulkley fit, $\tau=\tau_{y}+k\dot{\gamma}^{n}$, where $\tau_{y}=4.3\pm 0.3\text{Pa}$ (SD from best fit) is the dynamical yield stress. Other fit parameters are $k=0.55\pm 0.13$ and $n=0.85\pm 0.05$ (B) Storage $G^{\prime}$ and loss $G^{\prime\prime}$ moduli as a function of the deformation strain amplitude $\gamma$ for a oscillatory shear with angular frequency of $1\text{ rad/s}$.
@@
Fig. S8: Calibration steps for computing the correction function for the biovolume distribution. (A-B) Uncorrected concentration distribution of colonies measured in the cone-and-plate shear setup. Number of colonies counted during sampling: $N=10776$. (C-D) Reference concentration distribution measured in the inverted microscope after sedimentation of the colonies. Number of colonies counted during sampling, in panels C and D: $N=3066$ and $1455$, respectively. (E-F) Ratio between the reference and uncorrected concentration distributions Dashed lines indicate the fitted calibration parameters. (G-H) Normalized biovolume distribution measured in the cone-and-plate shear setup after the correction function. Error bars indicate the root mean squared of counting uncertainty and the size measurement uncertainty. Panels A,C,E,G display the distributions for small colonies ($l<5$), while panels B,D,F,H display the distributions for large colonies ($l>5$).